Status Report

NASA Space Station On-Orbit Status 11 Feb 2004

By SpaceRef Editor
February 11, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 11 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

FE Alexander Kaleri performed the (henceforth regular) daily 5-min. inspection of the newly installed BIO-5 Rasteniya-2/Lada-4 experiment and transferred the accumulated data files to a floppy disk, preparing them for downlink via Regul-Packet or OCA.   [Rasteniya-2 (“Plants-2”) researches growth & development of higher plants under spaceflight conditions in the Lada-4 greenhouse.  Experimental seeds of acacia-leaf type pea are planted between wicks in a root tray, with environmental control powered on.  Regular daily maintenance of the experiment involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, computer logging, and photo/video recording.]

Kaleri completed the installation of connections for the Russian “Matryoshka” radiation measurement experiment in the Service Module (SM).  Later in the day he connected to it the cables of the BITS2-12 onboard measurement telemetry system that collects engineering data from the setup.  Due to the large number of cables to be installed, Mike Foale was asked to assist Sasha if necessary.   [Matryoshka collects radiation measurements in the SM and DC-1 docking compartment as well as externally (after the EVA-9) for studies of on-orbit radiation and long-term dose accumulation, using six SPD dosimeters deployed throughout the RS as well as by in a spherical Matryoshka-R  “phantom”, which simulates a human torso.  Note: Matryoshka is the name for the traditional Russian set of nestling dolls.]

CDR Foale performed a portion of the scheduled HRF GASMAP (Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology) 30-day health check by bringing the unit up and letting it run on standby power for most of the day.  The ground downlinked its files during that time; then Mike turned everything off.   [A full-up 30-day health check will be scheduled next month.]

Earlier in the morning, Mike downlinked the FOOT (Foot/Ground Reaction Forces during Space Flight) experiment data of his most recent session (his second, of 1/15) through the HRF PC laptop.   [Due to scheduling conflicts, the PC was connected for the downlink to the rack for GASMAP, not to the HRF Workstation as usual.  This transfer of the FOOT data from the ADAS (Ambulatory Data Acquisition System) flash card to the HRF PC did not get all data files for FOOT or GASMAP, but is acceptable for the ground teams this one time.  For the 30-day health check, FOOT will again work through the HRF Workstation.]

Using the new procedure developed to prevent potential gas bubbles from getting into the Elektron-VM oxygen generator, Sasha Kaleri primed (filled) the KOV thermal loops’ EDV container with purified water from the multifiltration/purification column unit (BKO) while checking for any air bubbles.

The CDR retrieved the new U.S. TEPC (Tissue Equivalent Proportional Counter) spectrometer and its detector plus power, data and detector cables, all delivered on Progress 13P, and installed the radiation counter in the SM (panel #327) for activation and checkout.

Foale also terminated the 24-hour discharging process for the newly delivered spacesuit (EMU) batteries #2045 & #2046 in the Airlock (A/L)’s battery stowage assembly (BSA), then stowed them in the A/L after resetting their 50-day clock.

Mike had ~20 min. scheduled for bringing onboard hardcopy SODF (station operational data file) volumes up to date by placing new pages, launched on 13P, in the ISS Warning and ISS IFM Leak Kit Procedures books.

The CDR also performed the regular SOZh life support systems maintenance in the SM as well as the regular routine status checkup of autonomous Increment 8 payloads in the USOS.

After yesterday’s tightening of the four quick-disconnect (QD) couplings at the Lab window, early this morning Mike Foale’s task to perform the window depressurization procedure to return the inter-pane “Volume D” to vacuum was stricken off today’s schedule due to objections from Russian safety specialists.   [A new date for the procedure, which is intended to prevent flash condensation of any lingering humidity in the inter-pane space, has not been set yet.]

Mike Foale used the crewtime that became available due to the cancellation for the task-listed inspection of ELPS (emergency light power supply) units in the Node and Lab (except for forward Lab position.)

A recurring item on the Russian “job jar” task list is the continuing search, on the crew’s “free” time, for “missing” equipment, including some critical items required to support system operation that may be needed in the nearest future.  [Mike and Sasha were provided with a new list of this equipment, with pictures to facilitate search & identification.  The list calls out a BRPK condensate separation & pumping unit, two manually operated pumps, a current regulator, a hose, a BKV water treatment unit, and an IPD Draeger tubes cartridge belt.]

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Lagos, Nigeria (nadir pass over this burgeoning coastal city that is located on an offshore sand bar [like Galveston] and on the inland side of the lagoon), Snowy mountains, Pyrenees, Alps (Dynamic event.  Very clear conditions:  looking right for panoramas of both mountain ranges), Internal waves, Nicaragua (looking right towards the glint point to capture wave packets in the enigmatic zone on the Pacific side of central America.  Internal waves are known from this zone but the causes are unknown.  The wave patterns may indicate subsurface features), Internal waves, Central Pacific (looking right towards the glint disc as it passes across the western Tuamotu Island chain and internal waves forming in shallow inter-island zones), Phoenix, AZ (looking right towards this major desert urban region), Denver, CO (near-nadir pass:  One image may have been sufficient to capture the whole city), Palmyra Atoll (400mm-lens.  Nadir pass.  Detailed images were requested of the coral reefs), and Kingman Reef (400mm-lens.  Looking slightly left of track.  Detailed images were requested of the coral reefs).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:16pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off; SKV-2 is Off (repair incomplete).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.9; ppO2 (mmHg) — 155.1; ppCO2 (mmHg) — 2.4;
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.4.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 23.0.
  • Node:  Pressure (mmHg) — 747.07; temperature (deg C) — 23.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 749.46; temperature (deg C) — 25.1; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 749.56; temperature (deg C) — 24.1; shell heater temp (deg C) — 23.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.4
  • PMA-2:  Shell heater temp (deg C) — 10.1

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (non-suntracking, “night glider”/”sun slicer” drag reduction mode).
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); battery #5 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4115 kg (9072 lb) as of 2/5/04  [SM(755) + FGB(2701) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning,7:29am EST [= epoch]):

  • Mean altitude — 367.5 km
  • Apogee — 372.5 km
  • Perigee — 362.5 km
  • Period — 91.9 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0007411
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 85 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29844

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.